OpenStudy (lena772):

A solution of salt and water contains 50 grams of water per 100 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 25 milliliters of the solution. [Round off your answer to two decimal places]

OpenStudy (anonymous):

so for this you can set up a conversion table \[ \Big( \frac{25 \cancel{ml}}{} \Big)\Big( \frac{50 \ \cancel{g}}{100 \ \cancel{ml}} \Big)\Big( \frac{1 \ mole}{16 \cancel{g}} \Big)=\# \ \text{of moles of water}\] units of milliliters cancel, and units of grams cancel, leaving you with moles